OpenStudy (anonymous):

why 49−−√ is a rational number.

OpenStudy (anonymous):

Hmm, could you double-check that you've written out the numbers properly? :)

OpenStudy (anonymous):

4/9

OpenStudy (anonymous):

I'm not really sure what this bit means: \[49--\sqrt{}\]

OpenStudy (anonymous):

@BasketWeave I meant 4 over 9

OpenStudy (anonymous):

Ah ok, so just to check, the question is "Why is 4/9 a rational number?", right?

OpenStudy (anonymous):

yes @BasketWeave

OpenStudy (anonymous):

Or perhaps you meant this? \[\sqrt{\frac{ 4 }{ 9 }}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(4/9)^0.5=2/3 which is a rational number